# PyTomography: A Python Library for Medical Image Reconstruction

### v4

@@ Appendix A Validation of PyTomography
This section contains three examples used to validate PyTomography against other reconstruction software:
@@
    Monte Carlo PET Validation With CASToR: The ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image [25] used in the PET example was simulated in GATE without random, scatter, attenuation, or positron range modeling to obtain a ultra high count dataset consisting only of primary events. The data were reconstructed using LM TOF OSEM (2 iterations, 21 subsets) with both PyTomography and CASToR. Resulting reconstructions are shown in Figure 7. The mean activity ratio in grey to white matter was 2.05 (PyTomography) and 2.03 (CASToR). The noise in the image was estimated as the standard deviation of counts in white matter divided by the mean activity in white matter; the values were 0.329 (PyTomography) and 0.321 (CASToR). Apart from processing the ROOT files and building the sensitivity image, the time required for reconstruction in PyTomography was 14 s to build the system matrix and required 59 s/it for reconstruction. CASToR reconstruction was ran on a separate HP Z8 Computer Desktop with 2 Intel Xeon Gold 6136 CPUs at 3.00 GHz with a total of 48 CPU cores. The reconstruction was performed with 24 CPU cores and reconstruction required approximately 455 s per full iteration.

### v5

@@ Appendix A Validation of PyTomography
This section contains three examples used to validate PyTomography against other reconstruction software. Reconstruction times for cases 1 and 3 are shown in Table 3.
@@
    Monte Carlo PET Validation With CASToR: The ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image [28] used in the PET example was simulated in GATE without random, scatter, attenuation, or positron range modeling to obtain a ultra high count dataset consisting only of primary events. The data were reconstructed using LM TOF OSEM (2 iterations, 21 subsets) with both PyTomography and CASToR. Resulting reconstructions are shown in Figure 6. The mean activity ratio in grey to white matter was 2.05 (PyTomography) and 2.03 (CASToR). The noise in the image was estimated as the standard deviation of counts in white matter divided by the mean activity in white matter; the values were 0.329 (PyTomography) and 0.321 (CASToR). CASToR reconstruction was ran on a separate HP Z8 Computer Desktop with 2 Intel Xeon Gold 6136 CPUs at 3.00 GHz with a total of 48 CPU cores; this separate system was selected because it permitted use of more CPU cores in parallel during reconstruction, which is what CASToR is optimized for.